# Low-resolution Prior Equilibrium Network for CT Reconstruction

### v1

@@ 6. $\bullet$
    GRAD: gradient descent is used instead of the primal-dual algorithm to solve the low-level problem described in equation (9). It does not incorporate low-resolution image prior information. The network is a deep unrolled neural network with 10 iterations. The number of initialization channels is set to 5. The Xavier initialization and the mean squared loss of the reconstructed image and the ground truth are used in all experiments.
@@ 5.2 Datasets and settings
5.3 Evaluation Metrics
Two common metrics are adopted to evaluate the perceived quality of reconstructed images, including the peak signal-to-noise ratio (PSNR) and structural similarity (SSIM), which are defined as follows
$\mbox{PSNR}\left(\bm{{{x}}^{*}},\bm{x}\right)=10\times\log_{10}\frac{\left(\max\left(\bm{x}\right)-\min\left(\bm{x}\right)\right)^{2}}{\frac{1}{\mathrm{~{}N}}\left\|\bm{{{x}}^{*}}-\bm{x}\right\|^{2}},$
and
$\mbox{SSIM}\left(\bm{{{x}}^{*}},\bm{x}\right)=\frac{\left(2\mu_{\bm{{{x}}}^{*}}\mu_{\bm{x}}+c_{1}\right)+\left(2\sigma_{\bm{{{x}}^{*}}\bm{x}}+c_{2}\right)}{\left(\mu_{\bm{{{x}}}^{*}}^{2}+\mu_{\bm{x}}^{2}+c_{\mathrm{l}}\right)+\left(\sigma_{\bm{{{x}}}^{*}}^{2}+\sigma_{\bm{x}}^{2}+c_{2}\right)},$
with $\bm{{{x}}^{*}}$ and $\bm{x}$ as the ground truth image and the reconstructed image, respectively. The $\mu_{\bm{{{x}}}^{*}}$ and $\mu_{\bm{x}}$ represent the mean values of $\bm{{{x}}^{*}}$ and $\bm{x}$, $\sigma_{{\bm{x}}}$ and $\sigma_{\bm{x}}$ denote the standard deviations of $\bm{{{x}}^{*}}$ and $\bm{x}$, and $\sigma_{\bm{{{x}}^{*}}\bm{x}}$ represents the covariance between $\bm{{{x}}^{*}}$ and $\bm{x}$. Besides, $c_{1}$ and $c_{2}$ are small constants added to avoid division by zero.
(a) Results of the different iterations
@@ 5.3 Evaluation Metrics
Figure 5: Evaluation results of our LRPE with respect to different settings of parameters.
5.4 Test Settings and Parameter Choice
In this subsection, we present some numerical examples to evaluate the performance of our algorithm at different iterations. We conducted experiments with five different numbers of iterations, the PSNR of which are shown in Figure 5(a). As can be observed, as the number of iterations increases, the PSNR generally improves for all the limited-angle reconstruction problems. Since there is a trade-off between the reconstruction quality and computational complexity, we fix the total number of iterations to 10 to achieve a good balance. It is worth noting that by using a parameter-sharing strategy in the deep equilibrium model, the number of parameters is significantly reduced compared to the PD. Specifically, the number of parameters in our model is only 1/10th of the number of parameters in the PD. On the other hand, we analyze the effect of the hyper-parameter $\eta$ on the performance of limited-angle reconstruction. We performed a parameter sweep and obtained the results shown in Figure 5(b). From the results, we observed that the highest PSNR value was achieved when we set $\eta$ as $\eta=0.1$. Therefore, in all other experiments, we fixed the hyper-parameter $\eta$ to a value of 0.1.
Expanding the variable space is a common technique used in network optimization to improve stability during the training process. In our case, we expand the variable space by considering multiple variables $\bm{x}=[\bm{x}^{(1)},\bm{x}^{(2)},\ldots,\bm{x}^{(N_{p})}]$. In Table 1, we investigate the influence of different choices of $N_{p}$ on the reconstruction accuracy using limited-angle data. From the perspective of results obtained from various views in Table 1, we fix the value of $N_{p}$ to 5 for all experiments, which provides a good balance between reconstruction accuracy and computational efficiency.
@@ 5.4 Test Settings and Parameter Choice
Besides, we examine the impact of the number of epochs on the convergence of our network in the training process. The loss function, PSNR, and SSIM values are recorded and displayed in Figure 6. As can be seen, our LRPE network converges as the number of epochs increases. And we set the number of epochs to 28 for the subsequent experiments.
Figure 6: The loss function, PSNR, and SSIM values with respect to the numbers of epochs in our network. These curves were evaluated on the human phantom dataset with a scanning angular range of $150^{\circ}$ and 5% Gaussian noise.
@@ 5.5 Empirical data fidelity or learned data fidelity?
5.6 Sparse-view CT reconstruction
We conduct a further evaluation of the image quality of the proposed LRPE under low-dose scanning conditions, specifically sparse-view data. We utilize 60 views, 45 views, and 30 views projection data corrupted by 5% white Gaussian noises, where different reconstruction methods are evaluated in Table 4. Both PSNR and SSIM demonstrate that our LRPE has obvious advantages over other reconstruction methods. Figure 7 presents the reconstructed images obtained by the comparison methods on 30-view reconstruction. By comparing images (g) and (h), it is evident that the low-resolution image prior can significantly improve the reconstruction quality by preserving fine details and sharp edges. What’s more, the zoomed region exhibits that the deep equilibrium architecture can well maintain structural information.
@@ 5.6 Sparse-view CT reconstruction
Figure 7: The sparse-view reconstruction experiments were performed on the AAPM phantom dataset using 30 views and 5% Gaussian noise. The display window was set to [0, 1].
We further increased the noise level in the raw data to 10% white Gaussian noises and provided the quantitative results in Table 5. It can be observed that the FBP performs poorly in the presence of high-level noises, with a decrease in PSNR by 4 dB compared to previous experiments. In contrast, the learning-based methods are demonstrated to be less sensitive to noises. Similarly, our LRPE exhibits the best performance among all the deep learning-based algorithms.
@@
5.7 Limited-angle CT reconstruction
In this subsection, we evaluate the performance of our LRPE on the limited-angle reconstruction, where 5% Gaussian noises are corrupted into the projection data. Both PSNR and SSIM of the comparison methods are provided in Table 6. We can observe that the reconstruction qualities of all methods decrease as the scanning angle shrinks. Not surprisingly, our LRPE has the numerical advantage compared to other comparison algorithms, which provides a $0.8\mathrm{~{}dB}$ higher PSNR than the LRIP on $90^{\circ}$ limited-angle reconstruction task. Particularly, the projection data of the low-resolution prior used by LRIP is different from ours, which was computed using the downsampling matrix. Obviously, our setting is more reasonable and in accord with the CT scanner, which also gives reconstruction results with better qualities. Figure 8 presents the reconstruction results obtained by different methods for limited-angle reconstruction with a scanning range of $90^{\circ}$. The learning-based methods outperform both FBP and TV, which contain noticeable artifacts in regions within missing angles. Moreover, we can observe that our LRPE surpasses the SIPID, PD, FSR, and LRIP in preserving image details. Therefore, both quantitative and qualitative results confirm that the low-resolution image is a suitable prior for the ill-posed limited-angle reconstruction.
Table 6: Comparison of different methods on limited-angle data corrupted by $5\%$ Gaussian noises in terms of PSNR and SSIM, where time is recorded in ms.
@@ 5.7 Limited-angle CT reconstruction
In Figure 9, we present and compare the outputs of the LRIP, GRAD, and LEPE from the first iteration to the tenth iteration. As can be seen, as the number of iterations escalates, the PSNR value steadily rises, signifying the robust convergence of our method. In comparison to GRAD, it emphasizes the crucial significance of incorporating the low-resolution prior for incomplete reconstruction problems. On the other hand, the results of the LRIP in the first six iterations are not stable until the last two iterations provide meaningful results. Thus, the deep equilibrium scheme with convergence guarantee is a good choice for establishing the end-to-end reconstruction network.
@@
Last but not least, we conduct the comparison study on the running time of the evaluated methods, which are illustrated in Figure 10. Notably, excluding the running time of the TV method, it exhibits an execution time approximately ten times longer than that of the FBP. By analyzing the bar chart results, we can observe that although our method is not as rapid as the PD method, we incur only a 100-millisecond increase in running time to achieve superior reconstruction results.
@@ 6 Concluding Remarks
In this paper, we proposed a novel end-to-end equilibrium imaging geometric model for the ill-posed image reconstruction problems. The low-resolution image was used prior to guarantee that the proposed model can achieve ideal reconstruction results based on incomplete projection data. The advantages of our proposal were validated using a substantial amount of numerical experiments on sparse-view and limited-angle settings. The experimental results demonstrated that our model can enable accurate and efficient CT image reconstruction. Although our downsampling imaging geometric modeling was studied for the fan-beam imaging systems, it can be also used for other modern CT settings. Similar works include [53] and [52], where the low-resolution image prior was incorporated into the reconstruction network. Our method differs from [53] in that we utilized twice downsampled prior information, yet employed the same CT system, meaning we fixed identical $N_{views}$ and $N_{bins}$. In contrast, the prior information in [53] was obtained by hypothetically halving the receiver signal, which altered their $N_{bins}$. This necessitates a second scan of the patient, which is clinically prohibitive, whereas our approach obviates the need to rescan patients. On the other hand, Gao et al. [52] also requires two distinct sets of projection data, unlike our technique employs the same projection data derived from analyzing physical phenomena. Obviously, the low-resolution image obtained from the same scan can better maintain consistency with the high-resolution image, thereby preserving more details in the reconstructed images.

### v2

@@ 6. $\bullet$
    GRAD: gradient descent is used instead of the primal-dual algorithm to solve the low-level problem described in equation (9). It does not incorporate low-resolution image prior information. The network is a deep unrolled neural network with 10 stages. The number of initialization channels is set to 5. The Xavier initialization and the mean squared loss of the reconstructed image and the ground truth are used in all experiments.
@@ 5.2 Datasets and settings
Figure 6: The data error measured by L2 norm with respect to the numbers of epochs in our LRPE model. The curves represent the results for $150^{\circ}$ limited-angle (First row)and 60 sparse-view (Second row) with and without 5% Gaussian noises. On each row, the left one is the result without noises, and the right one is the result with 5% Gaussian noises.
5.3 Test Settings and Parameter Choice
In this subsection, we evaluate the choices of the parameters to the performance of our algorithm. There are several important parameters in Algorithm 1, including the number of stages $K$, the step size $\eta$ and the dimension of the variables in the network model.
We first study the influence of the stage number $K$ on the convergence of the algorithm. We use the squared error $\|\bm{u}^{K}-\bm{u}^{*}\|^{2}$ to measure the data error during the training process. Both the results of $150^{\circ}$ limited-angle and 60 sparse-view reconstruction with and without 5% Gaussian noises are presented in Figure 6. As we can see, our LRPE model demonstrates convergence as the number of epochs increases, regardless of whether there are 5, 10, or 15 stages (iterations). As Theorem 4.8 in [6] indicates, convergence to the optimal value can be achieved when the data itself is noise-free. If the observed data contains noises, the expected objective value of the fixed-point iteration converges to a neighborhood of the optimal value. We can observe from the numerical error of our LRPE method converges to lower error for cases without noises. When noises are introduced into observed data, the solution converges to much higher error, which is consistent with the theoretical analysis results in [6]. With the same number of iterations, more stages can optimize to a lower data error faster. Since there is a trade-off between the error reduction and training efficiency, we fix the total number of stages to 10 to achieve a good balance.
On the other hand, we also conducted experiments with different numbers of stages during the inference phase, where the LRPE models trained with 10 stages are used in the experiment. As can be observed in Figure 7(a), as the number of stages increases, the PSNR generally improves for all the limited-angle reconstruction problems. It is worth noting that by using a parameter-sharing strategy in the deep equilibrium model, the number of parameters is significantly reduced compared to the learned PD model. Specifically, the number of parameters in our model is only 1/10th of the number of parameters in the learned PD model. In what follows, we analyze the effect of the hyper-parameter $\eta$ on the performance of limited-angle reconstruction. We perform a parameter sweep and obtain the results shown in Figure 7(b), which models are used 10 stages. From the results, we observe that the highest PSNR value is achieved when we set $\eta$ as $\eta=0.1$. Therefore, in all other experiments, we fix the hyper-parameter $\eta$ to a value of 0.1.
(a) Results of the different stages
@@ 5.3 Test Settings and Parameter Choice
Figure 7: Evaluation results of our LRPE model with respect to different settings of parameters during inference process.
Expanding the variable space is a common technique used in network optimization to improve stability during the training process. In our case, we expand the variable space by considering multiple variables $\bm{u}=[\bm{u}^{(1)},\bm{u}^{(2)},\ldots,\bm{u}^{(N_{p})}]$. In Table 1, we investigate the influence of different choices of $N_{p}$ on the reconstruction accuracy using limited-angle data. From the perspective of results obtained from various views in Table 1, we fix the value of $N_{p}$ to 5 for all experiments, which provides a good balance between reconstruction accuracy and computational efficiency.
@@ 5.4 Empirical data fidelity or learned data fidelity?
Figure 8: The sparse-view reconstruction experiments are performed on the AAPM phantom dataset using 30 views and 5% Gaussian noise. The display window was set to [0, 1].
5.5 Experiments on the sparse-view reconstruction
We further evaluate the image qualities of our LRPE model under the low-dose sparse-view projection data. We utilize 60 views, 45 views, and 30 views projection data corrupted by 5% white Gaussian noises for evaluation, where the reconstruction results of different reconstruction methods are provided in Table 4. Both PSNR and SSIM demonstrate that our LRPE model outperforms other reconstruction methods. Figure 8 presents the reconstructed images obtained by the comparative methods on the 30-view reconstruction problem. By comparing images (g) and (h), it is evident that the low-resolution image prior can significantly improve the reconstruction quality by preserving fine details and sharp edges. Moreover, the zoomed region exhibits that the deep equilibrium architecture can well maintain structural information.
We further increase the noise level in the raw data to 10% white Gaussian noises and provide the quantitative results in Table 5. It can be observed that the FBP method performs poorly in the presence of high-level noises, with a decrease in PSNR by 4 dB compared to previous experiments. In contrast, the learning-based methods are demonstrated to be less sensitive to noises. Similarly, our LRPE model exhibits the best performance among all the deep learning-based algorithms.
@@ 5.5 Experiments on the sparse-view reconstruction
Table 6: Comparison of different methods on limited-angle data corrupted by $5\%$ Gaussian noises in terms of PSNR and SSIM.
@@
(a) FBP
(b) TV
(c) SIPID
(d) PD
(e) SFSR
(f) LRIP
(g) GRAD
(h) LRPE
Figure 9: Limited-angle reconstruction experiment with $150^{\circ}$ scanning angular range and 5% Gaussian noises.
5.6 Experiments on the limited-angle reconstruction
In this subsection, we evaluate the performance of our LRPE model on the limited-angle reconstruction, where 5% Gaussian noises are corrupted in the projection data. Both PSNR and SSIM of the comparison methods are provided in Table 6. We can observe that the reconstruction qualities of all methods decrease as the scanning angle shrinks. Not surprisingly, our LRPE model has the numerical advantage compared to other comparison algorithms, which provides a $0.8\mathrm{~{}dB}$ higher PSNR than the LRIP on $90^{\circ}$ limited-angle reconstruction task. Particularly, the projection data of the low-resolution prior used by LRIP model is different from ours, which was computed using the down-sampling matrix. Obviously, our setting is more reasonable and in accord with the CT scanner, which also gives better reconstruction results. Figure 9 and Figure 10 present the reconstruction results with a scanning range of $150^{\circ}$ limited-angle and $90^{\circ}$ limited-angle, respectively. As can be observed, the learning-based methods outperform both FBP and TV, which contain noticeable artifacts in regions within missing angles. Moreover, our LRPE model surpasses the SIPID, PD, FSR, and LRIP by providing more image details and continuous contours. Thus, both quantitative and qualitative results confirm that the low-resolution image is a suitable prior for the ill-posed limited-angle reconstruction. Furthermore, by comparing the reconstruction results of 150$\circ$ limited-angle and $90^{\circ}$, it is not difficult to find that the resolution of the reconstruction results of all methods significantly decreases with the loss of angles. Although our method can provide better structural information by introducing the low-resolution image prior, it still suffers details missing due to angle deficient. It is mainly because our low-resolution image is the direct reconstruction result of 90 limited-angle data. Thus, our future work includes to explore effective reconstruction methods for low-resolution images.
In Figure 11, we present and compare the output results of the LRIP, GRAD, and LRPE models during the iterative process. Although we have shown the numerical convergence of our LRPE model in the inference phase in Figure 7 (a), the visual comparison further demonstrates that our model can provide more meaningful intermediate results in the iterative process. Compared to GRAD model, our method can effectively improve the quality of reconstruction results by introducing a low-resolution image prior. On the other hand, the results of the LRIP model are not stable in the first six stages, and only provide meaningful results in the last two stages. Thus, the deep equilibrium model with convergence guarantee is a better choice for establishing a stable end-to-end reconstruction network.
@@ 5.6 Experiments on the limited-angle reconstruction
Last but not least, we conduct the comparison study on the running time of the evaluated methods, which are illustrate in Figure 12. Notably, excluding the running time of the TV method, it exhibits an execution time approximately ten times longer than that of the FBP. By analyzing the bar chart results, we can observe that although our method is not as rapid as the PD method, we incur only a 100-millisecond increase in running time to achieve superior reconstruction results.
@@ 6 Concluding Remarks
In this paper, we proposed a novel end-to-end equilibrium imaging geometric model for the ill-posed image reconstruction problems. The low-resolution image was used prior to guarantee that the proposed model can achieve good reconstruction results based on incomplete projection data. The advantages of our proposal were validated using a substantial amount of numerical experiments on sparse-view and limited-angle settings. The experimental results demonstrated that our model can enable accurate and efficient CT image reconstruction. Although our downsampling imaging geometric modeling was studied for the fan-beam imaging systems, it can be also used for other modern CT settings. Similar works include [29] and [17], where the low-resolution image prior was incorporated into the reconstruction network. The prior information in [29] was obtained by halving the receiver signal, requiring a second scan of the patient, which is clinically prohibitive. In contrast, our approach eliminates the need of re-scanning patients. On the other hand, Gao et al. [17] also requires two distinct sets of projection data. Obviously, the low-resolution image obtained from the same scan can better maintain consistency with the high-resolution image, thereby preserving more structural and fine information. According to the experimental results, although our LRPE model is superior to other algorithms in reconstruction results, there is still a problem of detail loss in the case of incomplete data sampling. Thus, enhancing the quality of low-resolution images through effective sinogram restoration methods is a possible way to improve the reconstruction results of our LRPE model on severe incomplete data problems. One of the future research directions is to enhance the resolution of reconstructed images under extreme limited-angle and low-dose scanning problems. In theoretical analysis, the noises in data present new challenges in estimating the gradients of network models. We plan to investigate the convergence of the deep equilibrium model based on stochastic gradient similar to [6].